OpenStudy (anonymous):

What is the minimum pressure in kPa that must be applied at 25 °C to obtain pure water by reverse osmosis from water that is 0.227 M in sodium chloride and 0.078 M in magnesium sulfate? Assume complete dissociation for electrolytes.

OpenStudy (anonymous):

Osmotic pressure=iCRT. Here you need to calculate iC seperately for two and add them up like i for NaCl=2. iC=0.454 And for MgSO4 it is 2*0.078=0.156. Hence iC=0.454+0.156=0.61. You know R and T So substitue and find the answer.